A researcher wants to create a 90% confidence interval for the population proportion, with the margin of error of 3%. She has no idea what the population proportion is and she cannot take a trial sample. Calculate the required sample size.
Answer #1

We have given that,

1-alpha=0.9

alpha=0.1

Margin error=3%=0.03

Z- value table=1.645

P=0.5

Formula,

Margin error=z*(√(p*(1-p)) /n)

0.03 =1.645*(√0.5*0.5/n)

0.03=1.645*0.5/√n

√n=1.645*0.5/0.03

=27.4167

n=751.68

n=752

Sample size is 752.